Consider a collection of envelopes consisting of

1 red envelope

1 blue envelope

3 green envelope3 and

1 yellow. envelope

If 2 envelopes are selected at​ random, without​ replacement, determine the probability that at least one is blue
Answer #1

Ans:

Total number of envelopes=1+1+3+1=6

Probability that at least one is blue=1-P(none is blue)

=1-(5/6)*(4/5)

=1-(4/6)

=1-(2/3)=1/3